Consider a 2230-lb automobile clocked by law-enforcement radar at a speed of 85.5 mph (miles/hour). If the position of the car is known to within 5.0 feet at the time of the measurement, what is the uncertainty in the velocity of the car?

DeltaV >= _____________ mph

If the speed limit is 75 mph, could the driver of the car reasonably evade a speeding ticket by invoking the Heisenberg uncertainty principle?

- YES

- NO
Answer #1

v   = 85.5 mph = 85.5 x 1609 m / 3600 s = 38.21 m/s

m = 2230lb = 1011.51 kg

x = 5 ft = 1.524 m

Heisenberg uncertainty principle,

deltaX * deltap =h' / 2

(1.524) * (m deltaV) = 6.634 x 10^-34 / 2*2*pi

deltaV = 3.425 x 10^-38 m/s = 7.66 x 10^-38 mph

there no way to get out. ANs. No